OpenStudy (anonymous):

Prove the identity. (1-cosx)(1+cosx)=1/1+cot^2(x)

OpenStudy (anonymous):

(1-cosx)(1+cosx) =1-cos^2(x) =sin^2(x) =??

OpenStudy (anonymous):

\(\Large\color{black}{Another~~~rule,~~~ \bf 1+cot^2(x)=sec^2(x) }\)

OpenStudy (anonymous):

\(\normalsize \color{black}{ \bf sin^2(x)=1/1+cot^2(x) }\) as you have worked out already, \(\normalsize \color{black}{ \bf sin^2(x)=1/sec^2(x) }\) \(\normalsize \color{black}{ \bf sin^2(x)=cos^2(x) }\) No.. IT is NOT an identity.....

OpenStudy (anonymous):

i understand it is this computer program we are using

OpenStudy (anonymous):

my bad, hold on

OpenStudy (anonymous):

thanks for your time

OpenStudy (anonymous):

\(\normalsize \color{black}{ \bf sin^2(x)=1/1+cot^2(x) }\) \(\normalsize \color{blue}{ Rule:~\bf 1+cot^2(x)=csc^2(x) }\) so it IS an identity, sorry.

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

you get the idea though... that's the main thing.

OpenStudy (anonymous):

sin^2(x)=1/csc^2(x)=1/1+cot^2(x)

OpenStudy (anonymous):

sweet

OpenStudy (anonymous):

yes.... :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you welcome.. I'll try not to confuse you again if we hopefully meet again.